OpenStudy (anonymous):

Manuel traveled 6 hours non stop to mexico, a total of 253 miles. He took a train part of the way, which averaged a speed of 50mph, and then took a bus the remaining distance, which averaged 40 mph. How long was manuel of the train??? Please help ! i am not good with word problems i need help setting it up!

OpenStudy (anonymous):

lets call the distance travelled by train \(x\) so the distance travelled by bus is \(253-x\)

OpenStudy (anonymous):

the time travelled by train is \(\frac{x}{50}\)and the time travelled by bus is \(\frac{253-x}{40}\)

OpenStudy (anonymous):

since time is distance divided by rate as the total time is \(6\) you can solve \[\frac{x}{50}+\frac{253-x}{40}=6\]

OpenStudy (anonymous):

multiply both sides by 200 to clear the fractions and solve for \(x\)

OpenStudy (anonymous):

so x=65???

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so he was on the train for 65hrs??

OpenStudy (anonymous):

wait hold the phone

OpenStudy (anonymous):

is the question really "How long was manuel of the train?"

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you are told that in the question!

OpenStudy (anonymous):

Manuel traveled 6 hours non stop to mexico evidently he was on the train for 6 hours

OpenStudy (anonymous):

it says he traveled a total time of 6 hrs but it doesnt say how long he was on each

OpenStudy (anonymous):

oh that is the question how long in each

OpenStudy (anonymous):

yes not a very good question

OpenStudy (anonymous):

ok so he was on the train for 65 miles

OpenStudy (anonymous):

yes that is what i got when i solved it

OpenStudy (anonymous):

the time he was on the train is \(\frac{65}{50}=1.3\) hours

OpenStudy (anonymous):

awesome thank you !!

OpenStudy (anonymous):

or one hour and twenty minutes

OpenStudy (anonymous):

yw